Your research team analysis the light of a mysterious object in space. By using a spectrometer,
you can observe the following spectrum of the object. The Hα line peak is clearly visible:
(a) Mark the first four spectral lines of hydrogen (Hα, Hβ, Hγ, Hδ) in the spectrum.
(b) Determine the radial velocity and the direction of the object’s movement.
(c) Calculate the distance to the observed object.
(d) What possible type of object is your team observing?
1
Expert's answer
2020-06-22T11:13:10-0400

(a) Mark the first four spectral lines of hydrogen:


(b) The rest wavelength of alpha-line is 656.28 nm, we have 800 nm, which gives the redshift of


"z=\frac{800}{656.28}-1=0.219"

The radial velocity is


"v=zc=65.65\cdot10^6\text{ m/s}."

The redshift (and velocity) is positive, the object is receding.

(c) Calculate the distance to the observed object by Hubble's law:


"r=\frac{v}{H_0}=\frac{65650\text{ km/s}}{66.93\text{ km/s/MPc}}=981\text{ MPc}."

(d) The possible type of object our team is observing is a cluster galaxy. It can probably be LEDA 25177.